sphenix-tpc-l AT lists.bnl.gov
Subject: Sphenix-tpc-l mailing list
List archive
[[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.31 at 9:30 AM ET
- From: "Hughes, Charles [PHYSA]" <chughes2 AT IASTATE.EDU>
- To: Thomas K Hemmick <Thomas.Hemmick AT stonybrook.edu>, David Baranyai <david.baranyai AT cern.ch>, Martin Purschke <purschke AT bnl.gov>, "Vasquez, Joel" <jvasquez AT bnl.gov>, Stephen Boose <boose AT bnl.gov>, Seth Howell <seth.howell AT stonybrook.edu>, "sphenix-tpc-l AT lists.bnl.gov" <sphenix-tpc-l AT lists.bnl.gov>, "Huang, Jin" <jhuang AT bnl.gov>, "divaldo95 AT gmail.com" <divaldo95 AT gmail.com>, Ross Corliss <ross.corliss AT stonybrook.edu>, Michael Lenz <mlenz AT bnl.gov>, "Kimelman, Benjamin" <benjamin.kimelman AT vanderbilt.edu>, Luke Legnosky <luke.legnosky AT stonybrook.edu>, "Haggerty, John" <haggerty AT bnl.gov>, "Mandracchia, Kevin" <kmandracchia AT bnl.gov>, Yeonju Go <ygo AT bnl.gov>, "Azmoun, Bob" <azmoun AT bnl.gov>, "Miraglia, Damon" <dmiraglia AT bnl.gov>, Gabor David <gabor.david AT stonybrook.edu>, "Sakaguchi, Takao" <takao AT bnl.gov>, "osbornjd91 AT gmail.com" <osbornjd91 AT gmail.com>, "Robert P. Pisani" <pisani AT bnl.gov>, Thomas K Hemmick <tkhemmick AT gmail.com>, Frank Toldo <fatoldo AT bnl.gov>, "Kuczewski, John" <jkuczewski AT bnl.gov>, shana Prifte <sprifte AT bnl.gov>, "josborn1 AT bnl.gov" <josborn1 AT bnl.gov>, Kristina Finnelli <kfinnelli AT bnl.gov>, "balazs.ujvari AT cern.ch" <balazs.ujvari AT cern.ch>, "Hughes, Charles [PHYSA]" <chughes2 AT IASTATE.EDU>, "Platte, Christopher W" <christopher.w.platte AT vanderbilt.edu>
- Subject: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.31 at 9:30 AM ET
- Date: Sun, 30 Mar 2025 17:00:00 +0000
Hello All,
Let's meet MONDAY 03.31 at 9:30 AM ET (3:30 PM CET) to discuss ongoing operations/analysis/planning for the TPC. In particular:
-
HV: @Yeonju Go has made analysis of the TPC gains without and with B field on for the HCAL singles data. She sees the gains as slightly lower WITH B field compared to WITHOUT (see below). One possible explanation for this is that both runs were taken on consecutive nights (03/24 B field OFF and 03/25 B field ON) which had rising gas pressure in the TPC with similar temperatures. This would mean <P/T> was greater on the night the B field was on compared to when it was off. It's possible the drop in gain due to weather may have been slightly more than the expected increase in gain due to magnet. We should discuss this. We should also discuss the weather correction, which still needs to be implemented. Finally, we can discuss HV operations this week with the anticipated beam and collisions (which may already be here).
AVG Ratio B Field ON/OFFSouth R30.97South R21.02South R10.82North R30.90North R20.88North R10.79
-
LV/FEEs: @Kuczewski, Johnuploaded the latest firmware on Saturday afternoon. We can discuss how this went plus anything we need to do on our end . Additionally, @Yeonju Go also observed a large increase in the R1 noise for the latest runs with magnet. I did see some wandering pedestals/stuck on stupid in OnlMon for the B field on data - NS 2 R1, NS 7 R1, SS 9 R1, SS 12 R1, SS 5 R1 - but these don't correspond to the modules @Yeonju Go highlights. If the general RMS noise is higher in the B field, that would be new to me. I took a pedestal run with the full B field on 03.25 and did not see any difference.
-
Gas/Cooling: Our Online and Offline Monitoring are strongly hinting at a change in the gas that started absorbing drift electrons sometime between 8:00 AM 03/27 and 6:30 PM 03/28. We see sloping pulse heights in the OnlMon and a z dependence of dE/dx in the offline - exactly as what we saw last year at the end of the 2-week AuAu run (though less severe at time of writing). @Robert P. Pisani has already done some forensics with no clear indication that these symptoms can be correlated to a gas change. We should discuss.
-
Line Laser: TPC group has been very hard at work making the line laser work. Last week we fired 9N and 6S. We learned the complete overhaul of the trigger system indeed results in the expected jitter < 50 ns. Much more work remains including aligning remaining benches and threading those into the TPC. We will have to discuss continuing operations in the coming week when the laser running is relegated to APEX and Maintenance days. We should also discuss the offline analysis of the data (which @Hughes, Charles [PHYSA] and @josborn1 AT bnl.gov have started to look at unpacking) which will be important. Finally, we can get any additional updates from @divaldo95 AT gmail.com and @Finnelli, Kristina.
-
Diffuse Laser: Last week we were able to resurrect the diffuse laser and prove all 6 were still working and send over a TPC only run to @Kimelman, Benjamin and @Platte, Christopher W for analysis. However, we are still not able to include the diffuse laser in the data taking because we cannot cap its rate at 100 Hz. Right now, its rate is still tied 1:1 to the GL1 (which means if the GL1 is firing at 4 kHz, the diffuse laser will trigger at 4 kHz - which it is not designed to do). We should discuss this with @Martin Purschke and @Huang, Jin so we can solve this problem.
-
Contingency Box: @Mandracchia, Kevin has been hard at work making the contingency boxes. We can hear any update from him.
-
Any other pressing business
Indico Link:https://indico.bnl.gov/event/27507/
-
[[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.31 at 9:30 AM ET,
Hughes, Charles [PHYSA], 03/30/2025
-
Re: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.31 at 9:30 AM ET,
John Haggerty, 03/31/2025
- Re: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.31 at 9:30 AM ET, Thomas K Hemmick, 03/31/2025
-
Re: [[Sphenix-tpc-l] ] TPC Meeting TOMORROW MONDAY 03.31 at 9:30 AM ET,
John Haggerty, 03/31/2025
Archive powered by MHonArc 2.6.24.